when the base of a ladder that had been resting flat against the wall is pulled out 4 meters the top of the ladder descends 1/5 the length of the ladder. What is the length of the ladder?
Answer by edjones(8007)

After the base of the ladder is pulled away 4m we have a right triangle.
Let L be the length of the ladder.
b=4, c=L, and a=L-.2L=.8L
a^2+b^2=c^2
(.8L)^2+4^2=L^2
.64L^2+16=L^2
.64L^2-L^2=-16
.36L^2=16
.6L=4
L=4/.6
=6 2/3 m